How do you compute algebraically (not by graphing) the intercepts of any function?

For example, what are the x-intercept(s) and y-intercept(s) of 2x – 5y = 8 without graphing the equation? Remember that intercepts are points and points are written as ordered pairs.

Answer by mananth(16946)

2x – 5y = 8

...
For y intercept plug x=0
0-5y=8
/-5
y=-8/5
The co-ordinates are (0,-8/5)
..
For x intercept
plug y = 0
2x+0=8
x= 4
(4,0) are the co-ordinates
..
m.ananth@hotmail.ca